OpenStudy (anonymous):

Write a formula for this sequence: 8, -1, -10, -19,... A. 8 - 9n B. 17 - 9n C. -19 - 9n

OpenStudy (anonymous):

so n is terms... for n1 we have ... 17-9*1=17-9=8 for n2 17-9*2=17-18=-1 for n3 17-9*3=17-27=-10 so we get... 8,-1,-10 ... exactly as we have in a row... so the formula should be at B i think xD

OpenStudy (anonymous):

thanks(: